OpenStudy (anonymous):

how can I Prove that it is true through an algebraic proof, identifying each step? (x+a)^2=x^2+2ax+a^2

OpenStudy (anonymous):

(x+a)(x+a)=?

OpenStudy (anonymous):

\[(x+a)^2=(x=a)(x+a)=x(x+a)+a(x+a)=x^2+ax+ax+a^2=x^2+2ax+a^2\]

OpenStudy (anonymous):

so when its broken down into "solving" it would be (x+a)^2=x^2+2ax+a^2 (x+a)(x+a)=x(x+a)+a(x+a) x^2+ax+ax+?

OpenStudy (anonymous):

the last term is a^2 ! :)

OpenStudy (anonymous):

thanks:) so that's the answer? or would I have to do more?